A student, starting from rest,slides down a water slide. On the way down, a kinetic frictionalforce (a nonconservative force) acts on her. Thestudent has a mass of 69.0 kg, and the height of the water slide is11.3 m. If the kinetic frictional force does -5.60 103 J of work, how fast is the student going atthe bottom of the slide?
Answer #1
Potential energy at the top = U = mgh
Energy at the bottom = K = Friction energy
--
U = mgh = (69kg)(9.8)(11.3m) = 7641.06
--
K = -5.60x10^3
--
U - K = 7641.06 - 5.60x10^3 = 2042 J
--
Solve for v:
--
1/2mv^2 = 2042
v = 7.69 m/s
--
--
Hope this helps.